Solve the following for x:
__x_=_2__
x-3 3

Answer by checkley71(8403)

x/(x-3)=2/3 cross multiplying we get:
3x=2(x-3)
3x=2x-6
3x-2x=-6
x=-6 answer.
proof
-6/(-6-3)=2/3
-6/-9=2/3
2/3=2/3